# Negative Feedback Training: A Novel Concept to Improve Robustness of NVCIM DNN Accelerators

### v3

@@ 1. Introduction
Deep Neural Networks (DNNs) have ushered in a profound transformation in our society. However, accelerating DNN inference, which involves substantial vector-matrix computational loads, is hindered by frequent data movement between memory and processing units. This von Neumann bottleneck (et al., 2016a) could be solved by Computing-In-Memory (CIM) DNN accelerators (et al., 2016c). Leveraging emerging Non-Volatile Memory (NVM) devices, NVCIM DNN accelerators perform parallel in-situ data processing, significantly outperforming their CMOS-based counterparts in terms of energy efficiency and density (et al., 2023d). However, the inherent non-idealities of NVM devices, such as device variation, contribute to the deviation in device conductance values (et al., 2011, 2015b, 2009). As a result, such deviations affect the represented model weights after programming and, consequently, have a detrimental impact on the inference accuracy of NVCIM DNN accelerators (et al., 2021b).
@@
Achieving reliable DNN inference with unreliable NVM substrates is challenging. There exists a variety of hardware and software solutions to address this, such as device upgrade (et al., 2015a), write-verify (et al., 2020d), training regularization (et al., 2023a), neural architecture search (et al., 2021b, 2020b), and Bayesian network (et al., 2021a). Among them, noise-injection (or noise-aware) training (et al., 2020a, 2022d, 2023c) has emerged as one of the most popular and effective methods with no need for accelerator hardware modification. Simply exposing DNNs to the Gaussian noise (for modeling device variation) (et al., 2020c, 2022b, 2022a) at the training stage, it enhances the model’s noise tolerance. However, state-of-the-art (SOTA) noise-injection training methods are subject to issues like limited accuracy improvement, less confident prediction (or increased epistemic uncertainty (et al., 2023b)), and hard-to-reach training convergence under large variations. Our observation is that this mainly stems from the mismatch between the non-deterministic noise and the deterministic training for the following reasons: First, only a limited number of noise samples are exposed to the network during limited training epochs, hindering the network from fully capturing the pattern of noise behavior. Second, compared with vanilla training, the random noise samples in noise-injection training offer more diverse optimization directions to the network, but some may lead to incorrect states, and the associated uncertainty exacerbates this trend. Therefore, the network may fail in converging to the optimal state during training, or the well-trained model may not generate high-quality predictions when confronted with variations in weight during inference.
@@ 2. Background and Related Work
NVM device non-ideal factors. NVM devices are susceptible to various sources of variations and noise, encompassing spatial and temporal variations. Spatial variations stem from defects that arise during fabrication, while temporal variations result from stochastic fluctuations in the device material. Unlike spatial variations, temporal variations are usually independent of the device but are subject to the programmed value (et al., 2018). For example, temporal variations can lead to different conductance values for the same NVM device when identical programming pulses are applied. In our research, we consider non-idealities to be uncorrelated among NVM devices, but our method can be adapted to account for other distributions of variations with appropriate modifications.
Prior Work. To address the DNN performance degradation due to above NVM device variations, two main research directions exist: reducing device conductance variation (hardware) and improving DNN robustness in the presence of device variation (software). For example, device improvements can be achieved through advancements in material development and fabrication processes (et al., 2015a), and write-verify (et al., 2020d) operations can be applied during device programming. Recent research indicates that selectively write-verifying only critical devices can improve the accuracy (et al., 2022c). Additionally, new circuits have been proposed to address device variations. On the software front, CorrectNet (et al., 2023a), which employs modified Lipschitz constant regularization during DNN training, is designed for robust DNNs. Other software-based methods involving training model weights with noises, designing architectures, and pruning techniques (et al., 2021b, a, 2020b) are also proposed. In this work, we focus on noise-injection training–one of the most important software-based solutions to this problem.
@@ 4.1. Model Weight Perturbation
We used parameter settings based on previous works (et al., 2020d, a, 2022c). Specifically, we set $K=2$, while the value of $M$ depended on the specific model configuration. In this study, we choose $M=8$, indicating 8-bit precision for a single DNN weight and 2-bit precision for a single device conductance. For device variation, we adopt a Gaussian distribution with $\Delta{g}\sim\mathcal{N}(0,\sigma_{d}^{2})$, where the $\sigma_{d}$ represents the relative standard deviation of conductance corresponding to the maximal conductance of a single device. We impose the constraint $\sigma_{d}\leq 0.4$ as this falls within a reasonable range achievable through device-level optimizations, such as write-verify based on measurement results. Our model and parameter settings align with those of RRAM devices described in (et al., 2020d).
@@ 4.2. Experimental Setup
We compare NFT with two baselines: 1) vanilla training (W/O Noise) and 2) Gaussian noise-injection training (W/ Noise) (et al., 2020a, 2022d). We did not assess NFT against other orthogonal methods like NAS-based DNN topology design or Bayesian Neural Networks, as NFT can be used in conjunction with them.
@@ 4.3. Effectiveness of Accuracy Improvement
For experiments, we use the VGG-8 (et al., 2014) backbone on MNIST and CIFAR10 datasets, and the ResNet-18 (et al., 2016b) backbone on MNIST, CIFAR10, CIFAR100, and Tiny ImageNet datasets.
@@ 4.4. Uncertainty and Convergence
Device variation amplifies epistemic uncertainty, leading to increased output uncertainty. We utilize Expected Kullback–Leibler (EKL) Divergence (et al., 2023b) (the lower the better) to quantify the impact of device variation on uncertainty. To ensure a fair comparison, accuracy is excluded from the analysis. Specifically, among all correct predictions in noisy inference, we calculate the Kullback–Leibler Divergence between each softmax output and its corresponding label. The results are the average EKL divergence for each correct prediction, shown in Figure 4. Compared to the W/O Noise baseline, W/ Noise baseline enhances accuracy but at the cost of increased uncertainty. In contrast, our NFT not only achieves even higher accuracy than the noise-injection training baseline but also maintains low uncertainty and high confidence in the output. In cases where device variation is too substantial for effective predictions using the vanilla training baseline, its EKL divergence appears slightly lower than that of NFT, as it generates completely random and meaningless predictions.
@@ 4.5. Ablation Study
Different $N$ of negative feedback: Table 1 shows the noisy inference accuracy when we try to use different numbers of negative feedback $Out_{n}$ in NFT. It shows that our setting achieves the best accuracy, which is $N=3$ for OVF and $N=4$ for IRS. Although continuing to increase the number of negative feedback tends to result in a slight improvement in accuracy, we maintained the current settings (marked as $*{}^{*}$) due to considerations of time and energy efficiency.
@@
Pretrained Models: To assess whether training time and energy can be saved by pretrained models, we use a pretrained model trained through the vanilla training method for 200 epochs (accuracy 92.71% with $\sigma_{d}=0$) in NFT. The noisy inference results are presented in Table 2. The findings indicate that the optimal approach for NFT is training from scratch, as implemented in our previous experiments (marked as $*{}^{*}$).

### v4

@@ 1. Introduction
Deep Neural Networks (DNNs) have ushered in a profound transformation in our society. However, accelerating DNN inference, which involves substantial vector-matrix computational loads, is hindered by frequent data movement between memory and processing units. This von Neumann bottleneck (Chen et al., 2016) could be solved by Computing-In-Memory (CIM) DNN accelerators (Shafiee et al., 2016). Leveraging emerging Non-Volatile Memory (NVM) devices, NVCIM DNN accelerators perform parallel in-situ data processing, significantly outperforming their CMOS-based counterparts in terms of energy efficiency and density (Zhang et al., 2023). However, the inherent non-idealities of NVM devices, such as device variation, contribute to the deviation in device conductance values (Rizzi et al., 2011; Raty et al., 2015; Chen et al., 2009). As a result, such deviations affect the represented model weights after programming and, consequently, have a detrimental impact on the inference accuracy of NVCIM DNN accelerators (Yan et al., 2021).
@@
Achieving reliable DNN inference with unreliable NVM substrates is challenging. There exists a variety of hardware and software solutions to address this, such as device upgrade (Degraeve et al., 2015), write-verify (Shim et al., 2020), training regularization (Eldebiky et al., 2023), neural architecture search (Yan et al., 2021; Jin et al., 2020), and Bayesian network (Gao et al., 2021). Among them, noise-injection (or noise-aware) training (et al., 2020; Yang et al., 2022; Yan et al., 2023) has emerged as one of the most popular and effective methods with no need for accelerator hardware modification. Simply exposing DNNs to the Gaussian noise (for modeling device variation) (Qin et al., 2020; Yan et al., 2022a; Wan et al., 2022) at the training stage, it enhances the model’s noise tolerance. However, state-of-the-art (SOTA) noise-injection training methods are subject to issues like limited accuracy improvement, less confident prediction (or increased epistemic uncertainty (Gawlikowski et al., 2023)), and hard-to-reach training convergence under large variations. Our observation is that this mainly stems from the mismatch between the non-deterministic noise and the deterministic training for the following reasons: First, only a limited number of noise samples are exposed to the network during limited training epochs, hindering the network from fully capturing the pattern of noise behavior. Second, compared with vanilla training, the random noise samples in noise-injection training offer more diverse optimization directions to the network, but some may lead to incorrect states, and the associated uncertainty exacerbates this trend. Therefore, the network may fail in converging to the optimal state during training, or the well-trained model may not generate high-quality predictions when confronted with variations in weight during inference.
@@ 2. Background and Related Work
NVM device non-ideal factors. NVM devices are susceptible to various sources of variations and noise, encompassing spatial and temporal variations. Spatial variations stem from defects that arise during fabrication, while temporal variations result from stochastic fluctuations in the device material. Unlike spatial variations, temporal variations are usually independent of the device but are subject to the programmed value (Feinberg et al., 2018). For example, temporal variations can lead to different conductance values for the same NVM device when identical programming pulses are applied. In our research, we consider non-idealities to be uncorrelated among NVM devices, but our method can be adapted to account for other distributions of variations with appropriate modifications.
Prior Work. To address the DNN performance degradation due to above NVM device variations, two main research directions exist: reducing device conductance variation (hardware) and improving DNN robustness in the presence of device variation (software). For example, device improvements can be achieved through advancements in material development and fabrication processes (Degraeve et al., 2015), and write-verify (Shim et al., 2020) operations can be applied during device programming. Recent research indicates that selectively write-verifying only critical devices can improve the accuracy (Yan et al., 2022b). Additionally, new circuits have been proposed to address device variations. On the software front, CorrectNet (Eldebiky et al., 2023), which employs modified Lipschitz constant regularization during DNN training, is designed for robust DNNs. Other software-based methods involving training model weights with noises, designing architectures, and pruning techniques (Yan et al., 2021; Gao et al., 2021; Jin et al., 2020) are also proposed. In this work, we focus on noise-injection training–one of the most important software-based solutions to this problem.
@@ 4.1. Model Weight Perturbation
We used parameter settings based on previous works (Shim et al., 2020; et al., 2020; Yan et al., 2022b). Specifically, we set $K=2$, while the value of $M$ depended on the specific model configuration. In this study, we choose $M=8$, indicating 8-bit precision for a single DNN weight and 2-bit precision for a single device conductance. For device variation, we adopt a Gaussian distribution with $\Delta{g}\sim\mathcal{N}(0,\sigma_{d}^{2})$, where the $\sigma_{d}$ represents the relative standard deviation of conductance corresponding to the maximal conductance of a single device. We impose the constraint $\sigma_{d}\leq 0.4$ as this falls within a reasonable range achievable through device-level optimizations, such as write-verify based on measurement results. Our model and parameter settings align with those of RRAM devices described in (Shim et al., 2020).
@@ 4.2. Experimental Setup
We compare NFT with two baselines: 1) vanilla training (W/O Noise) and 2) Gaussian noise-injection training (W/ Noise) (et al., 2020; Yang et al., 2022). We did not assess NFT against other orthogonal methods like NAS-based DNN topology design or Bayesian Neural Networks, as NFT can be used in conjunction with them.
@@ 4.3. Effectiveness of Accuracy Improvement
For experiments, we use the VGG-8 (Simonyan and Zisserman, 2014) backbone on MNIST and CIFAR10 datasets, and the ResNet-18 (He et al., 2016) backbone on MNIST, CIFAR10, CIFAR100, and Tiny ImageNet datasets.
@@ 4.4. Uncertainty and Convergence
Device variation amplifies epistemic uncertainty, leading to increased output uncertainty. We utilize Expected Kullback–Leibler (EKL) Divergence (Gawlikowski et al., 2023) (the lower the better) to quantify the impact of device variation on uncertainty. To ensure a fair comparison, accuracy is excluded from the analysis. Specifically, among all correct predictions in noisy inference, we calculate the Kullback–Leibler Divergence between each softmax output and its corresponding label. The results are the average EKL divergence for each correct prediction, shown in Figure 4. Compared to the W/O Noise baseline, W/ Noise baseline enhances accuracy but at the cost of increased uncertainty. In contrast, our NFT not only achieves even higher accuracy than the noise-injection training baseline but also maintains low uncertainty and high confidence in the output. In cases where device variation is too substantial for effective predictions using the vanilla training baseline, its EKL divergence appears slightly lower than that of NFT, as it generates completely random and meaningless predictions.
@@ 4.5. Ablation Study
Different $N$ of negative feedback: Table 1 shows the noisy inference accuracy when we try to use different numbers of negative feedback $Out_{n}$ in NFT. It shows that our setting achieves the best accuracy, which is $N=3$ for OVF and $N=4$ for IRS. Although continuing to increase the number of negative feedback tends to result in a slight improvement in accuracy, we maintained the current settings (marked as ) due to considerations of time and energy efficiency.
@@
Pretrained Models: To assess whether training time and energy can be saved by pretrained models, we use a pretrained model trained through the vanilla training method for 200 epochs (accuracy 92.71% with $\sigma_{d}=0$) in NFT. The noisy inference results are presented in Table 2. The findings indicate that the optimal approach for NFT is training from scratch, as implemented in our previous experiments (marked as ).